The potential GDP and actual GDP of an economy are $25000 and $20000 respectively. What is the GDP gap? If the natural rate of unemployment is 5%, what is the actual rate of unemployment?


1
Expert's answer
2021-02-07T17:59:23-0500

"\bold {Answers}"

GDP gap = $5,000 or 20%


Actual unemployment rate = 15%


"\bold {Solutions}"

GDP Gap "=" Potential GDP - Actual GDP

"= Y^* - Y"

"= \$25,000 - \$20,000"

"\bold {= \$5,000}"


GDP gap as a percentage of potential GDP "= \dfrac {Y^* - Y}{Y^*} × 100\%"


"= \dfrac {\$5,000}{\$25,000} × 100\%"


"\bold {= 20\%}"


According to Okun's Law, the relationship between output gap and unemployment rate is given by:

"\dfrac {Y^* - Y}{Y^*} = 2(u - u^*)" where "u^*" Is the natural rate of unemployment and "u" is the actual unemployment rare.

It is given that "u^* = 0.05"

Therefore,

"0.20 = 2(u - 0.05)"

Dividing by 2 both sides gives:

"0.10 = u - 0.05"

"=> u = 0.10 + 0.05"

"=> \bold {u = 0.15}"


Thus, the actual rate of unemployment, according to Okun's law is 15%.